OpenStudy (daenio):

If the force of gravity between a book of mass 0.50 kg and a calculator of 0.100 kg is 1.5 × 10-10 N, how far apart are they? Note: The distance found is between the centres of mass.

OpenStudy (anonymous):

Force = G m1 m2 / r^2

OpenStudy (daenio):

Do I have to convert the mass into g?

OpenStudy (anonymous):

nah. take them as it is (in kg)

OpenStudy (daenio):

\[1.5 x 10^{-11} N = (6.67 x 10^{-11} N*m ^{2}/kg ^{2}) (0.50 kg) (0.100kg) / r ^{2}\]

OpenStudy (daenio):

\[r ^{2}= \sqrt(6.67 x 10^{-11} N*m ^{2}/ kg ^{2} (0.50kg) (0.100kg) / 1.5 x 10^{-10} N\]

OpenStudy (daenio):

= 1.49 x 10^-11

OpenStudy (daenio):

\[=1.49 x 10^{-11} m\]

OpenStudy (daenio):

Is that correct?

OpenStudy (daenio):

@telltoamit

OpenStudy (daenio):

@maheshmeghwal9

OpenStudy (daenio):

@CarlosGP

OpenStudy (anonymous):

Your formula is ok:\[ r=\sqrt{Gm_1m_2/F}=\sqrt{6,67\times10^{-11}\times 0.5 \times0.1/1.5\times10^{-10}}=\sqrt {???}=???m\] But you made an error in the calculation, they are much, much closer

OpenStudy (anonymous):

Give your result and I'll check it

OpenStudy (daenio):

I have no idea.

OpenStudy (shane_b):

@CarlosGP: I think you mean much, much *farther* apart. @Daeno: The answer you gave was\[1.49 x 10^{-11} m\]So something went wrong in your calculation of the exponents. It should have been:\[0.149m\]

OpenStudy (anonymous):

I do not know what made me think of 14.9 cm shorter than 14.9 pico-meters. Things of the age, surely. Thanks for the correction

OpenStudy (shane_b):

Happens to me all the time :)

OpenStudy (anonymous):

Equations+Exponents+English as Second Language=Fatal Errors :)

OpenStudy (shane_b):

lol...I don't even have the second language excuse!

OpenStudy (daenio):

I don't know where I went wrong with the equation.

OpenStudy (daenio):

Can someone please clarify on what I should've done?

OpenStudy (maheshmeghwal9):

ya answer is 0.149 meters. as @Shane_B is correct 100%. u made calculation mistake @Daenio :/ just improve that:)

OpenStudy (daenio):

Can you break down the calculation because I don't know how to get 0.149m.